### v1

@@ Abstract
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. However, achieving these guarantees in practice often requires stronger conditions than the null hypothesis of primary interest. For instance, sign-change tests for mean zero require symmetry and fail to control finite-sample error for non-symmetric mean-zero distributions. We investigate whether such limitations stem from specific test choices or reflect a fundamental inability to construct valid randomization tests for certain hypotheses. We develop a framework providing a simple necessary and sufficient condition for when null hypotheses admit randomization tests. Applying this framework to one-sample tests, we provide characterizations of which nulls satisfy this condition for both finite and continuous supports. In doing so, we prove that certain null hypotheses — including mean zero — do not admit randomization tests. We further show that nulls that admit randomization tests based on linear group actions correspond only to subsets of symmetric or normal distributions. Overall, our findings affirm that practitioners are not inadvertently incurring additional Type 1 error when using existing tests and further motivate focusing on the asymptotic validity of randomization tests.
@@ 1 Introduction
Consider the problem of testing a null given a finite sample of data. If this null satisfies a group invariance property referred to as the “randomization hypothesis,” then one can construct randomization tests that obtain exact finite-sample Type 1 error control. Intuitively, this property requires that, under the null, there are transformations of the data that preserve the distribution of the data. In part due to the allure of finite-sample error control, there is a large methodological literature on randomization tests and a similarly large applied literature that makes use of these tests (for a review, see Ritzwoller et al. 2024).
@@
Randomization tests that are used in practice often require stronger conditions than the null hypothesis of primary interest. This is a problem because it re-introduces an inability to control Type 1 error: we may over-reject when the data is drawn from a distribution that satisfies the hypothesis of primary interest but not the stronger conditions. For instance, a commonly-used randomization test for the null of mean zero given i.i.d. observations is constructed by repeatedly recomputing a test statistic under all possible sign changes on the observations. However, this test only properly controls finite-sample error when the data is drawn from a symmetric distribution, and thus may over-reject when data is drawn from a non-symmetric distribution with mean zero.
@@
This paper investigates whether the failure to control Type 1 error for certain null hypotheses (e.g. the null of mean zero) is due to the specific randomization tests used in practice (e.g. the use of sign changes) or due to a more general inability to construct randomization tests that are valid for the hypotheses. To formalize this distinction, we develop a framework to characterize which null hypotheses can and cannot be tested using a randomization test. In doing so, we provide a simple necessary and sufficient condition for a null hypothesis to satisfy the randomization hypothesis that avoids group-theoretic properties that typically accompany the theory of randomization tests.
We then apply our framework to analyze one-sample randomization tests. We show that for certain null hypotheses — such as the null of mean zero — there does not exist a randomization test that properly controls finite-sample error. These results are related to results concerning the non-existence of reasonable inference procedures of certain non-parametric hypotheses (Bahadur & Savage 1956, Lehmann & Loh 1990, Romano 2004). By focusing on randomization tests, we are able to develop conditions for the (non-)existence of finite-sample randomization tests for many null hypotheses and among different families of distributions. For instance, this finding affirms that existing applications of randomization tests are not inadvertently paying “too much” of a Type 1 error when using tests based on sign changes.
We also use our framework to inform researchers on how to look for and construct randomization tests for null hypotheses for which standard constructions are not appropriate. When the data is drawn from a finite support, we provide an explicit characterization of which null hypotheses admit a randomization test. When the data is instead drawn from a continuous support, we consider which null hypotheses admit a randomization test based on groups of linear actions. To the best of our knowledge, existing randomization tests are all based on such groups. We show that the set of one-sample null hypotheses that admit such a test are extremely limited: they only correspond to subsets of symmetric distributions and of normal distributions.
@@
As part of identifying whether a null hypotheses admits a randomization test, we also show how to use our framework to explicitly construct randomization tests. Throughout the paper, we provide explicit examples — including a construction of a normality test that makes use of rotation symmetries — that may be of theoretic or applied interest.
These conclusions yield a clear positive implication that applied work that uses randomization tests should continue to do so in the current manner. Our largely negative results also affirm and motivate a line of work that considers the asymptotic properties of randomization tests (see, e.g., Romano (1989, 1990), DiCiccio & Romano (2017), Canay et al. (2017), Lei & Bickel (2021), Pouliot (2024), Bai et al. (2024)). These works develop asymptotic properties of specific tests (such as the above test of a mean of zero) or of randomization tests more generally. Finally, our results also suggest that one-sample randomization tests that are not testing for symmetry or normality must necessarily come from non-linear transformations.
@@
The remainder of the paper is organized as follows. Section 2 reviews the randomization test procedure and develops our general framework. Section 3 applies our framework to the case of one-samples tests considering i.i.d. data supported on a finite set. Section 4 then considers the case where the data is supported on the real line. Section 5 concludes.
@@ 2.1 Construction of the randomization test
The construction follows Lehmann et al. (1986); see Zhang & Zhao (2023) for a recent formulation and discussion of the test. Let $X=(X_{1},\dots,X_{n})$ be data taking values in a sample space $\mathcal{X}$ with distribution $P\in\Omega$. Consider the problem of testing $H_{0}:P\in\Omega_{0}$, and suppose that there is a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\in P\in\Omega_{0}$. For ease of exposition, we develop the procedure when $\mathbb{G}$ is a finite group of size $M$.
Let $T:\mathcal{X}\to\mathbb{R}$ be any test statistic used to test the null. For every $x\in\mathcal{X}$, let $T^{(1)}(x;\mathbb{G})\leq\dots\leq T^{(M)}(x;\mathbb{G})$ be the ordered values of $T(gx)$ as $g$ varies in $\mathbb{G}$. Given nominal level $\alpha\in(0,1)$, let $k\equiv M-\lfloor M\alpha\rfloor,$ where $\lfloor M\alpha\rfloor$ denotes the largest integer less than or equal to $M\alpha$. Let $M^{+}(x)$ and $M^{0}(x)$ be the number of values $T^{(j)}(x)$ that are greater than $T^{(k)}(x)$ and equal to $T^{(k)}(x)$, respectively. Set
@@
$\phi(x;\mathbb{G})=\begin{cases}1&\text{if $T(x)>T^{(k)}(x;\mathbb{G})$}\\ a(x)&\text{if $T(x)=T^{(k)}(x;\mathbb{G})$}\\ 0&\text{if $T(x)<T^{(k)}(x;\mathbb{G})$}\end{cases},$ (1)
@@
Let $\phi$ be the randomization test as described above, and suppose that for any $g\in\mathbb{G}$, $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\in P\in\Omega_{0}$. Then $\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\alpha$ $\forall\;P\in\Omega_{0}$.
@@ 2.2 Framework
The set inclusions in (3) enforce that $\Omega_{0}$ is invariant to $\mathbb{G}$ while also ensuring that $\mathbb{G}$ is not invariant to all distributions in $\Omega$. The latter condition ensures that the test has power by excluding trivial choices of groups (such as the identity group) that are invariant to all distributions in $\Omega$.
Our analysis will focus on characterizing which null hypotheses do and do not satisfy the randomization hypothesis. For those that do, we will be interested in determining a group that allows for an implementation of the randomization test procedure. The following result provides a simple tool to answer both of these questions in a way that avoids group-theoretic concepts.
@@
where $\Omega_{f}\equiv\big\{P\in\Omega\;:\;fX\stackrel{{\scriptstyle d}}{{=}}X,X\sim P\big\}.$ When this holds, a randomization test can be constructed by using the group generated by $f$ or any collection of functions that satisfy (4).
@@
Proposition 1 shows that to study which null hypotheses satisfy the randomization hypothesis, it suffices to consider individual bijective functions. Given a bijective function that satisfies (4), we can then implement the randomization test procedure by taking the group generated by the function. We will later illustrate this approach with specific examples.
In the rest of the paper, we will apply these results to analyze one-sample tests. With some recycling of notation, we will accordingly consider $X=(X_{1},\dots,X_{n})$ with $X_{i}$ drawn i.i.d. from distribution $P\in\Omega$ supported on a (subset of) the real line. In this notation, it is useful to let $\Omega$ denote the set of distributions for $X_{i}$.
@@ 3 One-sample tests on a finite support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the finite set $\mathcal{A}\equiv\{\alpha_{1},\dots,\alpha_{K}\}\subset\mathbb{R}$. We can then represent $P$ by the associated function $p:\mathcal{A}\to[0,1]$ and represent $\Omega_{\text{discrete}}$ via
@@
By Proposition 1, a null hypothesis $\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there is some function $f:\mathcal{A}^{n}\to\mathcal{A}^{n}$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$. On a finite support, we can explicitly write $\Omega_{f}$ as
$\Omega_{f}=\bigcap_{x\in\mathcal{A}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(f_{j}(x))\right\}.$ (6)
@@
Second, Proposition 2 highlights that the null hypotheses that satisfy the randomization hypothesis are only those that place equal mass on different sets of points (accounting for multiplicities). As such, it is natural to expect that null hypotheses that are not characterized by such features likely do not satisfy the randomization hypotheses. In what follows, we use Proposition 2 to show that null hypotheses about $k$-th moments (including the mean) and quantiles do not satisfy the randomization hypothesis. These results and our explicit proofs illustrate how a researcher can use our results to determine if a null hypothesis of interest does or does not satisfy the randomization hypothesis.
@@
The proofs of Propositions 3 and 4 follow by showing that whenever we have $x,y\in\mathcal{A}^{n}$ with the two vectors not being equal as sets, then we can construct an explicit distribution in the null that does not place equal mass on $x$ and $y$. The conclusions then follow by Proposition 2.
@@
The cardinality requirements in Propositions 3 and 4 are necessary to avoid edge cases. For example, when $|\mathcal{A}|=3,$ then for any set $\mathcal{A}=\{-L,0,L\}$, the null of mean zero is equivalent to the null of symmetric distributions, which satisfies the randomization hypothesis using any group of sign changes. The proof of Proposition 3 provides an edge case when $|\mathcal{A}|=4$. The proof of Proposition 4 reveals that the statement holds when $|\mathcal{A}|=2$ as long as we exclude $p=1/2$. When $|\mathcal{A}|=2$ and $p=1/2$, we can construct the null hypothesis by using any group of sign changes.
@@ 4 One-sample tests on a continuous support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the real line. In what follows, we first consider groups of linear transformations, and explicitly characterize the set of all null hypotheses that are invariant to any such group. We then consider arbitrary groups, and develop and apply a necessary condition for a null hypothesis to satisfy the randomization hypothesis.
@@ 4.1 Groups of linear transformations
A group of linear transformations $\mathbb{G}$ is one in which each group element $g$ corresponds to a matrix $A_{g}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The following result shows that if $\mathbb{G}$ is a group of linear transformations, then the set of distributions that is invariant to $\mathbb{G}$ is either a subset of the symmetric distributions with mean zero or of the Gaussian distributions.
@@
Let $P\in\Omega_{\text{cts}}$, where $\Omega_{\text{cts}}$ is a subset of distributions supported on the real line. Let $\mathbb{G}$ be a group of linear transformations. Then exactly one of the following is true: 1. $\Omega_{\mathbb{G}}=\Omega_{\text{cts}}$, 2. $\Omega_{\mathbb{G}}=\{P\in\Omega_{\text{cts}}:\text{$P$ symmetric around $0$}\}$, 3. $\Omega_{\mathbb{G}}\subseteq\{N(\mu,\sigma^{2})\in\Omega_{\text{cts}}:\mu\in\mathbb{R},\sigma^{2}\in\mathbb{R}_{>0}\}$, or 4. $\Omega_{\mathbb{G}}=\emptyset$.
The proof follows by representing $gX$ via $AX$ for some matrix $A$. If the matrix is diagonal, then the only possible invariance is over the symmetric distributions. If the matrix is not diagonal, then $AX$ and $X$ are random vectors of i.i.d. random variables with at least one component of $AX$ featuring a non-trivial linear combination of multiple components of $X$. By the Darmois-Skitovich Theorem, this implies $X_{i}$ must be i.i.d. Gaussian.
Proposition 5 establishes that if we restrict to groups formed by linear maps, one-sample randomization tests only generalize permutation tests — which are invariant to all distributions in one-sample settings — in their ability to test if a distribution belongs to the symmetric distributions with mean zero or to a subset of the Gaussian distributions. The former can be implemented using the group of sign changes. Example 2 of Supplementary Appendix A shows how to construct randomization tests of the latter.

### v2

@@ Abstract
Randomization tests deliver exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. In practice, achieving these guarantees often requires stronger conditions than the null hypothesis of primary interest. For example, sign-change tests of mean zero require symmetry and need not control finite-sample size for non-symmetric mean-zero distributions. We investigate whether the mismatch between the null and the invariance conditions required for exactness reflects the use of particular transformations or a more fundamental limitation. We provide a simple necessary and sufficient condition for a null hypothesis to admit a randomization test. Applying this framework to one-sample problems, we characterize the nulls that admit randomization tests on finite supports and derive impossibility results on continuous supports. In particular, we show that several common nulls, including mean zero, do not admit randomization tests. We further show that, among one-sample tests using linear group actions, the admissible nulls are limited to subsets of symmetric or Gaussian distributions. These results confirm that the absence of exact finite-sample validity is inherent for many commonly studied nulls and that practitioners using existing tests are not foregoing feasible exact alternatives.
@@ 1 Introduction
Consider the problem of testing a null hypothesis given a finite sample of data. If this null satisfies a group invariance property referred to as the “randomization hypothesis,” then one can construct randomization tests that obtain exact finite-sample Type 1 error control. Intuitively, this property requires that, under the null, certain transformations of the data leave its distribution unchanged. In part due to the appeal of finite-sample validity, there is a large methodological literature on randomization tests and a similarly large applied literature using these methods (for a review, see Ritzwoller et al. 2024).
@@
In practice, randomization tests often rely on invariance conditions that are stronger than the null hypothesis of primary interest. This is a problem because it re-introduces an inability to control Type 1 error: the test can over-reject when the data is drawn from a distribution that satisfies the hypothesis of primary interest but not the stronger conditions. For example, a commonly used randomization test for the null of mean zero with i.i.d. observations is based on sign changes. This test controls finite-sample size only under symmetry and may over-reject for non-symmetric mean-zero distributions.
@@
This paper investigates whether the failure to control Type 1 error for certain null hypotheses (e.g. the null of mean zero) is due to the specific randomization tests used in practice (e.g. the use of sign changes) or from a more fundamental limitation. To formalize this distinction, we develop a framework that characterizes which null hypotheses admit randomization tests. We provide a simple necessary and sufficient condition for a null to satisfy the randomization hypothesis that avoids group-theoretic structure.
We then apply this framework to one-sample problems. We show that certain null hypotheses—such as the null of mean zero—do not admit randomization tests that achieve exact finite-sample validity. These results relate to classical non-existence results for nonparametric inference (Bahadur and Savage 1956, Lehmann and Loh 1990, Romano 2004). By focusing on randomization tests, we derive conditions for the existence and non-existence of exact finite-sample procedures across a range of null hypotheses and distributional classes. These results imply that existing applications of randomization tests, such as sign-change tests, are not overlooking alternative procedures that would achieve exact finite-sample validity.
We also provide guidance for constructing randomization tests when standard approaches are not applicable. On finite supports, we give an explicit characterization of nulls that admit randomization tests. On continuous supports, we study tests based on linear group actions. We show that, in one-sample settings, such tests are very limited: the admissible nulls correspond to subsets of symmetric or Gaussian distributions.
@@
As part of identifying whether a null admits a randomization test, we also show how our framework can be used to construct these tests. We do so by providing explicit examples — including a construction of a normality test that makes use of rotation symmetries — that may be of theoretic or applied interest.
Taken together, these results yield a positive implication for applied work: the use of standard randomization tests does not overlook alternative procedures that achieve exact finite-sample validity for many commonly studied nulls. They also affirm and motivate a focus on asymptotic properties of randomization tests (see, e.g., Romano (1989, 1990), DiCiccio and Romano (2017), Canay et al. (2017), Lei and Bickel (2021), Pouliot (2024), Bai et al. (2024)). Finally, our results suggest that one-sample randomization tests that go beyond symmetry or normality must rely on non-linear transformations.
@@
The remainder of the paper is organized as follows. Section 2 reviews the randomization test procedure and develops our framework. Section 3 studies one-sample tests with finite support. Section 4 considers continuous supports. Section 5 concludes.
@@ 2.1 Construction of the randomization test
The construction follows Lehmann et al. (1986); see Zhang and Zhao (2023) for a recent formulation. Let $X=(X_{1},\dots,X_{n})$ be data taking values in $\mathcal{X}$ with distribution $P\in\Omega$. Consider the problem of testing $H_{0}:P\in\Omega_{0}$, and suppose that there is a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\sim P\in\Omega_{0}$. For ease of exposition, we assume $\mathbb{G}$ is finite with size $M$.
Let $T:\mathcal{X}\to\mathbb{R}$ be any test statistic used to test the null. For every $x\in\mathcal{X}$, let $T^{(1)}(x;\mathbb{G})\leq\dots\leq T^{(M)}(x;\mathbb{G})$ be the ordered values of $T(gx)$ as $g$ ranges over $\mathbb{G}$. Given nominal level $\alpha\in(0,1)$, let $k\equiv M-\lfloor M\alpha\rfloor,$ where $\lfloor M\alpha\rfloor$ denotes the largest integer less than or equal to $M\alpha$. Let $M^{+}(x)$ and $M^{0}(x)$ be the number of values $T^{(j)}(x)$ that are greater than $T^{(k)}(x)$ and equal to $T^{(k)}(x)$, respectively. Set
@@
$\phi(x;\mathbb{G})\equiv\begin{cases}1&\text{if $T(x)>T^{(k)}(x;\mathbb{G})$}\\ a(x)&\text{if $T(x)=T^{(k)}(x;\mathbb{G})$}\\ 0&\text{if $T(x)<T^{(k)}(x;\mathbb{G})$}\end{cases},$ (1)
@@
Suppose that for any $g\in\mathbb{G}$, $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\sim P\in\Omega_{0}$. Then $\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\alpha$ $\forall\;P\in\Omega_{0}$.
@@ 2.2 Framework
The set inclusions in (3) enforce that $\Omega_{0}$ is invariant to $\mathbb{G}$ while also ensuring that $\mathbb{G}$ is not invariant to all distributions in $\Omega$. The latter condition rules out trivial groups (e.g. the identity) that are invariant to all distributions.
Our analysis will focus on characterizing which null hypotheses do and do not satisfy the randomization hypothesis. For those that do, we will be interested in constructing a group to implement the test. The following result provides a tool to answer both of these questions in a way that avoids group-theoretic concepts.
@@
where $\Omega_{f}\equiv\big\{P\in\Omega\;:\;fX\stackrel{{\scriptstyle d}}{{=}}X,X\sim P\big\}.$ When this holds, a randomization test can be constructed by using the group generated by $f$.
@@
Proposition 1 shows that to study which null hypotheses satisfy the randomization hypothesis, it suffices to consider individual bijective functions. Given such a function satisfying (4), a randomization test can be implemented using the group it generates.
In the remainder of the paper, we study one-sample problems with $X=(X_{1},\dots,X_{n})$ with $X_{i}$ i.i.d. from $P\in\Omega$, supported on a subset of the real line. Here, $\Omega$ denotes the set of distributions for $X_{i}$.
@@ 3 One-sample tests on a finite support
In this section, we suppose $X_{i}$ are drawn i.i.d. from $P$ supported on a finite set $\mathcal{A}\equiv\{\alpha_{1},\dots,\alpha_{K}\}\subset\mathbb{R}$. We represent $P$ by its probability mass function $p:\mathcal{A}\to[0,1]$ and define
@@
By Proposition 1, a null hypothesis $\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there exists a bijection $f:\mathcal{A}^{n}\to\mathcal{A}^{n}$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$. On a finite support, we can explicitly write $\Omega_{f}$ as
$\Omega_{f}=\bigcap_{x\in\mathcal{A}^{n}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(f_{j}(x))\right\}.$ (6)
@@
Second, Proposition 2 shows that null hypotheses satisfying the randomization hypothesis are characterized by equal probability mass assigned to distinct configurations (accounting for multiplicities). Null hypotheses not defined by such restrictions therefore do not satisfy the randomization hypothesis. In what follows, we apply Proposition 2 to show that null hypotheses defined by $k$-th moments (including the mean) and quantiles do not satisfy the randomization hypothesis. These results illustrate how the proposition can be used to determine whether a null hypothesis admits a randomization test.
@@
The proofs of Propositions 3 and 4 proceed by showing that for any $x,y\in\mathcal{A}^{n}$ that are not equal up to permutation, one can construct a distribution in the null that assigns different probabilities to $x$ and $y$. The result then follows from Proposition 2.
@@
The cardinality requirements in Propositions 3 and 4 exclude edge cases. For example, when $|\mathcal{A}|=3$ and $\mathcal{A}=\{-L,0,L\}$, the null of mean zero coincides with symmetry, which satisfies the randomization hypothesis via sign changes. The proof of Proposition 3 provides an analogous edge case for $|\mathcal{A}|=4$. For Proposition 4, the result continues to hold when $|\mathcal{A}|=2$ except at $p=1/2$, where sign changes yield a valid randomization test.
@@ 4 One-sample tests on a continuous support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the real line. We first characterize all null hypotheses that are invariant to linear group actions. We then consider arbitrary groups and develop a necessary condition for a null hypothesis to satisfy the randomization hypothesis.
@@ 4.1 Groups of linear transformations
A group of linear transformations $\mathbb{G}$ is one in which each element $g$ corresponds to a linear map $A_{g}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The following result characterizes the distributions invariant to such groups.
@@
Let $P\in\Omega_{\text{cts}}$, where $\Omega_{\text{cts}}$ is a subset of distributions supported on the real line. Let $\mathbb{G}$ be a group of linear transformations acting on $\mathbb{R}^{n}$. Then exactly one of the following is true: 1. $\Omega_{\mathbb{G}}=\Omega_{\text{cts}}$, 2. $\Omega_{\mathbb{G}}=\{P\in\Omega_{\text{cts}}:\text{$P$ symmetric about $0$}\}$, 3. $\Omega_{\mathbb{G}}\subseteq\{N(\mu,\sigma^{2})\in\Omega_{\text{cts}}:\mu\in\mathbb{R},\sigma^{2}>0\}$, or 4. $\Omega_{\mathbb{G}}=\emptyset$.
The proof proceeds by writing $gX=AX$ for some matrix $A$. If $A$ is diagonal, invariance requires each coordinate to be symmetric about zero. If $A$ is not diagonal, then at least one component of $AX$ is a non-trivial linear combination of multiple coordinates of $X$. Since both $AX$ and $X$ consist of i.i.d. components, the Darmois-Skitovich theorem implies that $X_{i}$ must be Gaussian.
Proposition 5 shows that, when restricted to linear group actions, one-sample randomization tests extend permutation-based invariance in limited ways. In particular, they can test membership in the class of distributions symmetric about zero or in subsets of the Gaussian family. The former can be implemented using sign changes. Example 2 in Supplementary Appendix A illustrates the latter.